Suppose that price of AAA stock at the end of next year depends on the state of the economy. The economy can have three states with the supplied probabilities and prices. What is your expected return if you purchase AAA today for $100? Also assume that AAA will pay a dividend of $10 at the end of the year. Enter your answer as a percent. Round your final answer to 2 decimals.

What is the standard deviation of AAA’s returns if you purchase AAA today for $100?

Expansion

Stable

Recession

Probability

.25

.5

.25

Price in one year

$130

$100

$80
Answer #1

Expected return=0.25*((130+10)/100-1)+0.50*((100+10)/100-1)+0.25*((80+10)/100-1)=12.50000%

Standard deviation=sqrt(0.25*((130+10)/100-1-12.50000%)^2+0.50*((100+10)/100-1-12.50000%)^2+0.25*((80+10)/100-1-12.50000%)^2)=17.85357%